OpenStudy (shreenithaa):

help!! turn in to an inequality Find the three smallest consecutive even integers whose sum is greater than 126.

OpenStudy (shreenithaa):

does anyone know this?

OpenStudy (shreenithaa):

@jennyrlz

OpenStudy (johnweldon1993):

Let (x) = an even integer, (x + 2) = the next even integer and (x+4) be the next even one So since the sum of these is greater than 126 we have \[\large x+(x+2)+(x+4) > 126\]

OpenStudy (shreenithaa):

i dont get it?

OpenStudy (johnweldon1993):

You dont get what I have written? Well Lets say I have an even number...lets call it x Now to get to the next CONSECUTIVE even number....I need to add 2 to the one I have....if I add 1...it becomes odd So now I have x .... and then x+2 And I need the NEXT one as well..so follow suite...add 2 to the previous number we have x.....then x+2.....and now x+2 + 2 = x + 4 So the 3 consecutive even numbers we have to be summed are \[\large x + (x+2) + (x+4)\] Since the question then states, this sum is greater than 126...we just set our number equation accordingly \[\large x + (x+2) + (x+4) > 126\]

OpenStudy (johnweldon1993):

Now we just combine like terms....and evaluate for 'x'

OpenStudy (shreenithaa):

oooh now i get it thnx